# Quantum cohomology of projective bundles

### v3

@@ 5.7. Semisimplicity
Setting $z=0$, this shows that the map sending $\phi_{i}p^{k}$ to $\bigoplus_{j=0}^{r-1}\partial_{\hat{\tau}^{i,k}}\varsigma_{j}$ gives a ring isomorphism $(H^{*}(\mathbb{P}(V)),\star_{\hat{\tau}})\cong\bigoplus_{j=0}^{r-1}(H^{*}(B),\star_{\varsigma_{j}(\hat{\tau})})$ over $\mathbb{C}(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$. This proves Corollary 1.8.
@@
We can describe a decomposition of the eigenvalues of the quantum multiplication by the Euler vector field of $\mathbb{P}(V)$ along $Q=\hat{\tau}=0$. Because $z\nabla_{z\partial_{z}}=E_{\mathbb{P}(V)}\star_{\hat{\tau}}+O(z)$ for the quantum connection $\nabla$ of $\mathbb{P}(V)$, Theorem 5.1 implies that the eigenvalues of $E_{\mathbb{P}(V)}\star_{\hat{\tau}}$ are the union of the eigenvalues of $E_{B}\star_{\varsigma_{j}(\hat{\tau})}$, $j=0,\dots,r-1$. When $Q=\hat{\tau}=0$, the quantum product $\star_{\varsigma_{j}(\hat{\tau})}$ for $B$ equals the cup product and the eigenvalues of $E_{B}\star_{\varsigma_{j}(\hat{\tau})}$ are all equal to the $H^{0}$ component $r\lambda_{j}=re^{2\pi\sqrt{-1}j/r}q^{1/r}$ of $E_{B}$ along $Q=\hat{\tau}=0$ (see Theorem 5.1(4)). See Figure 2 for a typical decomposition of eigenvalues. Similar pictures have appeared in [28, Figure 16], [34, Figure 1].

### v4

@@ 5.7. Semisimplicity
[…]
@@
We can describe a decomposition of the eigenvalues of the quantum multiplication by the Euler vector field of $\mathbb{P}(V)$ along $Q=\hat{\tau}=0$. Because $z\nabla_{z\partial_{z}}=E_{\mathbb{P}(V)}\star_{\hat{\tau}}+O(z)$ for the quantum connection $\nabla$ of $\mathbb{P}(V)$, Theorem 5.1 implies that the eigenvalues of $E_{\mathbb{P}(V)}\star_{\hat{\tau}}$ are the union of the eigenvalues of $E_{B}\star_{\varsigma_{j}(\hat{\tau})}$, $j=0,\dots,r-1$. When $Q=\hat{\tau}=0$, the quantum product $\star_{\varsigma_{j}(\hat{\tau})}$ for $B$ equals the cup product and the eigenvalues of $E_{B}\star_{\varsigma_{j}(\hat{\tau})}$ are all equal to the $H^{0}$ component $r\lambda_{j}=re^{2\pi\sqrt{-1}j/r}q^{1/r}$ of $E_{B}$ along $Q=\hat{\tau}=0$ (see Theorem 5.1(4)). See Figure 2 for a typical decomposition of eigenvalues. Similar pictures have appeared in [28, Figure 16], [38, Figure 1].
5.8. Reconstruction method of Hinault-Yu-Zhang-Zhang
[…]
We describe the initial conditions for $\Phi$ and $\varsigma_{j}$ along $Q=\hat{\tau}=0$. Recall that the isomorphism $\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}$ arises from the restriction of the isomorphism $\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}$ in (5.8) to the subspace $\mathsf{H}=\bigoplus_{k=0}^{r-1}H^{*}(B)\lambda^{k}$. Since $\hat{\tau}(\tau)|_{Q=\tau=0}=0$ by Theorem 4.2(7), we have $\Phi_{j}|_{Q=\hat{\tau}=0}=\Pi_{j}\circ\operatorname{FT}^{-1}|_{Q=\tau=0}$. We know by Theorem 4.2(8) that $\operatorname{FT}^{-1}(\phi_{i}p^{k})|_{Q=\tau=0}=\phi_{i}\lambda^{k}$. Therefore we have
$\Phi_{j}(\phi_{i}p^{k})|_{Q=\hat{\tau}=0}=\Pi_{j}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=e^{-\sigma_{j}|_{Q=\tau=0}/z}q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(\phi_{i}\lambda^{k})$
by Proposition 5.8. Consider the case $\phi_{i}p^{k}=1$. Since $q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(1)$ is of the form $\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}e^{2\pi\sqrt{-1}jc_{1}(V)/(rz)}(1+O(q^{-1/r}))$, its logarithm with respect to the cup product structure on $H^{*}(B)$ is well-defined. Because $\Pi_{j}(1)$ contains no negative powers of $z$, the initial value $\varsigma_{j}^{\circ}:=\varsigma_{j}(\hat{\tau})|_{Q=\hat{\tau}=0}=r\lambda_{j}+\sigma_{j}|_{Q=\tau=0}$ for $\varsigma_{j}$ can be calculated as
(5.11) $\varsigma_{j}^{\circ}=r\lambda_{j}+[z^{-1}]\log\left(q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(1)\right)$
where $z^{-1}](\cdots)$ denotes the coefficient of $z^{-1}$. The initial value $\Phi_{j}^{\circ}:=\Phi_{j}|_{Q=\hat{\tau}=0}$ for $\Phi_{j}$ is then given by
(5.12) $\Phi_{j}^{\circ}(\phi_{i}p^{k})=e^{-(\varsigma_{j}^{\circ}-r\lambda_{j})/z}q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(\phi_{i}\lambda^{k})$
for $0\leq i\leq s$ and $0\leq k\leq r-1$. Note that $\mathscr{F}_{j}(\phi_{i}\lambda_{k})$ can be calculated explicitly in terms of the Chern classes of $V$ via the procedure in §5.3.
We briefly describe the reconstruction of $\Phi$ and $\varsigma_{j}$ from their initial conditions (5.11), (5.12) by means of Birkhoff factorization. We leave the detailed verification to the reader, as the argument is completely parallel to the case of blowups treated in [29, §5.8]. We set:
(5.13) $s_{j}(\hat{\tau})=\varsigma_{j}(\hat{\tau})-\varsigma_{j}^{\circ}.$
[…]
$M=\bigoplus_{j=0}^{r-1}e^{-\varsigma_{j}^{\circ}/z}M_{B}(\varsigma_{j}^{\circ}+s_{j};Qq^{-c_{1}(V)/r}).$
This satisfies $M|_{Q=s_{0}=\dots=s_{r-1}=0}=\operatorname{id}$ and $M=\operatorname{id}+O(z^{-1})$. Then the composition $(\Phi^{\circ})^{-1}\circ M$ admits a unique Birkhoff facorization of the form
$(\Phi^{\circ})^{-1}\circ M=M^{\prime}\circ\Phi^{-1}$
where $M^{\prime}=\operatorname{id}+O(z^{-1})$ is a power series in $z^{-1}$ and $\Phi^{-1}$ is a power series in $z$ such that
$M^{\prime}$ $\in\operatorname{End}(H^{*}(\mathbb{P}(V)))\otimes\mathbb{C}[z^{-1}](\!(q^{-1/r^{\prime}})\!)[\![Q,s_{0},\dots,s_{r-1}]\!],$
$\Phi^{-1}$ $\in\operatorname{Hom}(H^{*}(B)^{\oplus r},H^{*}(\mathbb{P}(V)))\otimes\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,s_{0},\dots,s_{r-1}]\!],$
[…]